OpenStudy (anonymous):

I have a question regarding exercise 14 in chapter 4.4 of Strang's book. It is clear from the theory part that the missing element in the 2x2 matrix should be equal to q1^T*b. But when I try to work backwards through Gram-Schmidt, i don't get this answer. Probably something very silly, but i cannot figure out what i am doing wrong. To get back to vector b, i need to take a certain combinations of the columns of Q. That is: b = ?*q1+||B||*q_2. This gives b = ?*q1+B. From the first step of Gram-Schmidt we have b = B + (a^Tb)/(a^Ta)a, therefore ? ends up being equal (a^T*b).

OpenStudy (anonymous):

You've made a tiny error: From the first step of Gram-Schmidt we have \[b = B + \frac{A^Tb}{A^TA}A\] that's right. But if you plug it back in your equation \[ \begin{align*} b &= ?q_1+ B\\ B + \frac{A^Tb}{A^TA}A &= ?q_1+ B\\ \frac{A^Tb}{A^TA}A &= ?q_1\\ A^Tb\frac{A}{||A||^2} &= ?\frac{A}{||A||} \end{align*} \] you can see that \(?\) has to be \(\frac{A^Tb}{||A||} = \frac{A}{||A||}^Tb\) which is \(q_1^Tb\).

OpenStudy (anonymous):

Thanks! I know exactly what got me confused. The capital letters that are used to express vectors. \[A ^{T}A\] read as a matrix product to me. Damn this notation! :-)